### v1

# Electron-phonon coupling in magnetic materials using the local spin density approximation

Abstract
Magnetic materials are crucial for manipulating electron spin and magnetic fields, enabling applications in data storage, spintronics, charge transport, and energy conversion, while also providing insight into fundamental quantum phenomena. In numerous applications, the interaction between electrons and lattice vibrations, known as electron-phonon coupling, can be of significant importance. In that regard, we extend the EPW package to be able to interpolate the electron-phonon matrix elements combining perturbation theory and maximally localized Wannier functions. This allows to use dense momentum grids at a reasonable computational cost when computing electron-phonon-related quantities and physical properties. We validate our implementation considering ferromagnetic iron and nickel, where we explore the absence of phonon-driven superconductivity, finding that superconductivity is intrinsically suppressed. Furthermore, we evaluate the carrier resistivity at finite temperatures for both systems, considering the role of the magnetic phase in carrier transport. Our findings indicate that in the case of Fe, the primary contributor to resistivity is electron-phonon scattering. In contrast, for Ni, electron-phonon scattering constitutes less than one-third of the resistivity, underscoring a fundamental difference in the transport properties of the two systems.
I Introduction
Magnetic materials play an important role for technological applications, including magnetic cores in electric motors and generators, or magnetic memory devices in computers and mobile phones. From a historical standpoint, iron is arguably the most significant magnetic material. Iron can be alloyed to produce metals, insulators [1], superconductors [2], strong ferromagnetism [3], antiferromagnetism [4], thermal shielding at low and high temperatures and has good mechanical properties [5].
Interest in magnetic materials is at an all-time high due the end of Moore’s law [6] and other limitations that current electronic devices are facing. In that regard, magnetic materials are seen as an appealing solution leveraging the spin of the electron giving rise to the field of spintronics. Spintronic devices exploit both the electronic and spin degrees of freedom [7, 8, 9], offering promising logic and memory architectures with inherently low power consumption [10, 11, 12, 13]. In addition, spintronic devices are widely used for memory applications [14, 15, 16, 17]. Iron and nickel are ubiquitous in steels, alloys, and functional magnetic devices used in the energy sector. Thus, there is renewed effort to understand, search and characterize potential materials relevant for the mentioned applications [18, 19, 20, 21]. Apart from these advances in devices for classical computers, the field of quantum computing might also require magnetic compounds that present specific transport properties [22, 23], and certain robustness to thermal noise to build and shield the quantum bits [24, 25].
Given that the electrical and thermal conductivity in metals and semiconductors is strongly influenced by electron–phonon scattering, developing accurate first-principles methods to capture this mechanism is central to understanding and reducing energy losses in conductive materials. Although electron-phonon interactions have been studied for decades, accurately describing them has proven to be a challenging task. The situation becomes even more complicated when one considers the effect that phonons can have on the local magnetic structure [26, 27, 28, 29, 30], where in some cases they can even induce a non-zero transient magnetization in nonmagnetic materials [31, 32]. As a result, it is common to rely on effective models that take input from experiments [33, 34, 35], or previously computed ab initio quantities [36, 37, 38, 39]. Today, first-principles electron-phonon interactions are widely available thanks to density functional theory (DFT) [40, 41] and density functional perturbation theory (DFPT) [42, 43, 44, 45]. However, this approach is not free from challenges that can make actual calculations impractical. One of the main bottlenecks is the high computational cost required to sample the Brillouin zone in both electronic $\mathbf{k}$ and phononic $\mathbf{q}$ momentum space when integrating the different observables. In that regard, one of the most widespread solutions is to combine DFPT with maximally localized Wannier functions [46, 47] which enables interpolation at arbitrary momentum $\mathbf{k}$ and $\mathbf{q}$ of the electron-phonon matrix elements (EPME) $g_{mn\nu}(\mathbf{k},\mathbf{q})$, electronic eigenvalues $\varepsilon_{n\mathbf{k}}$ and eigenvectors $\psi_{n\mathbf{k}}$. Most implementations to date assume time-reversal symmetry. This assumption is only suited to study non-magnetic materials, hindering a global understanding of electron-phonon interactions. In fact, this is an important barrier to understanding the role of phonons in spintronics, caloritronics, or superconductivity and charge density waves in the presence of magnetism. More importantly, both electron and phonon dispersions can be affected by including the spin degree of freedom in the calculations while retaining the same exchange and correlation functional [48]. For example, neglecting the spin of iron leads to unstable phonons [49].
This limitation has been addressed using finite differences [48, 50, 51, 52, 53, 54, 55] but is limited due to the computational cost of supercells. Recently, support for DFPT with non-collinear magnetism and ultrasoft pseudopotentials has been implemented in Abinit [56, 57, 58, 49, 59] and Quantum ESPRESSO [60, 61, 62]. However, it has never been extended to the interpolation of EPMEs, which hinders realistic applications in magnetic materials. We present a first-principles implementation for electron–phonon properties in magnetic systems, validated and applied to Fe and Ni. Our results reveal the essential role of the magnetic phase in resistivity, the microscopic origin of scattering in different ferromagnets, and the absence of phonon-mediated superconductivity in both materials. These insights provide a foundation for designing magnetic materials with tailored transport properties for energy applications. In this work, we extend the capabilities of the EPW code [46, 63, 64] to support collinear magnetism, unlocking the path to study electron-phonon related properties in magnetic materials with high accuracy and at reasonable computational cost. To that end, we exploit the recent developments in the Abinit code on spin-dependent electron-phonon interactions [57] to validate our implementation. This choice is guided by the documented small discrepancy between Abinit and Quantum ESPRESSO [65, 66, 67]. Although collinear magnetism serves as an approximation to its non-collinear counterpart when considering spin-orbit coupling (SOC), its computational efficiency is markedly superior, while retaining a good degree of accuracy.
We study the carrier transport properties of ferromagnetic (FM) body-centered cubic (BCC) Fe and face-centered cubic (FCC) Ni when electron phonon interactions are accounted. Our findings indicate that the omission of magnetism in Fe leads to the emergence of imaginary (soft) phonon modes, confirming previous works. As sometimes done in the literature, excluding imaginary phonon modes overestimated resistivity due to the large contribution of modes on the onset of being imaginary. In contrast, Ni is a weaker ferromagnet with similar ferromagnetic and non-magnetic phonon band dispersions but a very different change of resistivity with temperature. Overall, the inclusion of magnetism reduces resistivity by a factor of three and seven for Fe and Ni, respectively. Finally, we compute the phonon-driven superconducting transition temperature ($T_{\rm c}$) in Fe and Ni and find vanishingly small values, which shows that neither material would be superconducting, even if their Cooper pairs could survive in the FM magnetization field. This work opens the gate towards accounting for the role of phonons in magnetic systems, and possible superconducting states in magnetic materials (such as iron arsenides and possibly cuprates), which are highly relevant for new energy-efficient technologies and energy loss in conducting materials. This work is organized as follows: we first present the theory in Sect. II, details about our computational framework in Sect. III, the interpolation of basic electron, phonon and electron-phonon quantities in Sect. IV, the calculation of physical properties and electron-phonon properties in Sect. V, and the final remarks and conclusions in Sect. VI.
II Theory
We start with the definition of the EPME in the phonon mode representation [68]
[…]
where $\omega_{\mathbf{q}\nu}$ is the frequency of the phonon mode $\nu$ with quasi-momentum $\mathbf{q}$, $e_{\kappa\alpha\nu\mathbf{q}}$ the associated eigenvector, $M_{\kappa}$ is the mass of the $\kappa$-th atom in the unit cell $\mathbf{R}_{p}$, and $\partial_{\mathbf{q}\kappa\alpha}V$ is the variation of the effective potential with respect to the perturbation involving atom $\kappa$ displaced in Cartesian direction $\alpha$. This definition has to be updated when considering non-collinear magnetism as
[…]
where $\sigma$ and $\sigma^{\prime}$ are the spinor indices. From the definition of $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$, we can identify two terms. The first is proportional to $\omega_{\mathbf{q}\nu}^{-\frac{1}{2}}$. The second is the spin-dependent deformation potential $D^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$ which can be computed as [69]
$D^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})=\sqrt{\frac{m_{\mathrm{u}}\omega_{\mathbf{q}\nu}}{\hbar}}\Big|g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})\Big|,$ (3)
with $m_{\mathrm{u}}$ being the atomic mass constant. $D^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$ gives a measure of how much a phonon perturbation has changed the internal self-consistent potential felt by the electrons. In taking the absolute value of $|g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})|$, we take the root mean square average over any degenerate manifolds in $m$, $n$ and $\nu$. Note that the definitions of $D^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$ and $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$ are also valid for the case of nonmagnetic compounds with and without SOC. The main property exploited in the nonmagnetic case are the following extra symmetries in the spin indices $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})]^{*}=(-1)^{\sigma+\sigma^{\prime}}g^{\bar{\sigma}^{\prime}\bar{\sigma}}_{mn\nu}(\mathbf{-k},\mathbf{-q})$, with $\bar{\sigma}$ being the opposite spin to $\sigma$. This means that if $\sigma=\sigma^{\prime}$ then $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})]^{*}=g^{\bar{\sigma}^{\prime}\bar{\sigma}}_{mn\nu}(\mathbf{-k},\mathbf{-q})$, and if $\sigma\neq\sigma^{\prime}$ then $g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})]^{*}=-g^{\bar{\sigma}^{\prime}\bar{\sigma}}_{mn\nu}(\mathbf{-k},\mathbf{-q})$ (e.g. for heavy elements with spin orbit coupling but no magnetization). These two results are a consequence of time-reversal symmetry, which is absent in magnetic compounds.
In practice, performing magnetic calculations with SOC, commonly known as non-collinear magnetic calculations, is computationally expensive (typically by a factor of $2^{3}=8$ for cubic scaling DFT) in systems with several atoms. The collinear approximation consists in separating the two spin channels into up ($\uparrow$) and down ($\downarrow$) without explicit interaction between them, thereby only doubling the cost of the calculations, regardless of the system size. This allows us to define a spin-resolved EPME:
$g^{\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})\equiv\delta_{\sigma^{\prime}\sigma}g^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})$ (4)
and it follows that the deformation potential becomes
$D^{\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})=\delta_{\sigma^{\prime}\sigma}D^{\sigma^{\prime}\sigma}_{mn\nu}(\mathbf{k},\mathbf{q}).$ (5)
In systems with many valence electrons, we are commonly interested in the deformation potential of several bands. In that regard, for each spin channel we can consider
$D^{\sigma}_{\nu}(\mathbf{k},\mathbf{q})=\sqrt{\sum_{mn\in\mathcal{M}}[D^{\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})]^{2}}$ (6)
as the deformation potential of interest, with the sum over the electronic bands $m,n$ restricted to the bands in the Wannier manifold $\mathcal{M}$. We also extend the mode- and momentum-resolved electron-phonon coupling $\lambda_{\mathbf{q}\nu}$, which in the double delta approximation at the collinear level becomes [70]
$\lambda_{\mathbf{q}\nu}\equiv\sum_{\sigma}\lambda_{\mathbf{q}\nu}^{\sigma}=\sum_{\sigma mn\mathbf{k}}\frac{w_{\mathbf{k}}}{N(\varepsilon_{\rm F})\omega_{\mathbf{q}\nu}}|g_{mn\nu}^{\sigma}(\mathbf{k},\mathbf{q})|^{2}\\ \times\delta(\varepsilon_{\rm F}-\varepsilon^{\sigma}_{n\mathbf{k}})\delta(\varepsilon_{\rm F}-\varepsilon^{\sigma}_{m\mathbf{k+q}}),$ (7)
where $N(\varepsilon_{\textrm{F}})$ is the total density of states at the Fermi energy $\varepsilon_{\rm F}$ and $w_{\mathbf{k}}$ are the $\mathbf{k}$-point weights normalized to one. We note that another definition exists for the spin-resolved electron-phonon coupling using the partial density of states at the Fermi level $N^{\sigma}(\varepsilon_{\rm F})$ [58]. However, this convention will suffer from divergences in half- and semi-metals, where only one spin channel contributes to the total density of states. The present version should be used systematically for spin polarized cases, heeding the factor of 2 between $N^{\sigma}$ and $N$ in the unpolarized case.
Another important quantity is the mode and momentum resolved phonon linewidth $\gamma_{\mathbf{q}\nu}$
$\gamma_{\mathbf{q}\nu}\equiv\sum_{\sigma}\gamma^{\sigma}_{\mathbf{q}\nu}=2\pi\omega_{\mathbf{q}\nu}\sum_{\sigma\mathbf{k}mn}w_{\mathbf{k}}|g^{\sigma}_{mn\nu}(\mathbf{k},\mathbf{q})|^{2}\\ \times\delta(\varepsilon_{\rm F}-\varepsilon^{\sigma}_{n\mathbf{k}})\delta(\varepsilon_{\rm F}-\varepsilon^{\sigma}_{m\mathbf{k+q}}),$ (8)
which can be used to compute the Eliashberg spectral function
$\alpha^{2}F(\omega)=\frac{1}{2\pi N(\varepsilon_{\rm F})}\sum_{\sigma\mathbf{q}\nu}w_{\mathbf{q}}\frac{\gamma^{\sigma}_{\mathbf{q}\nu}}{\omega_{\mathbf{q}\nu}}\delta(\omega-\omega_{\mathbf{q}\nu}),$ (9)
where $w_{\mathbf{q}}$ are the $\mathbf{q}$-point weights normalized to one. The Eliashberg spectral function is used to determine the contribution of specific phonon modes to the total electron-phonon coupling strength defined as
$\lambda\equiv\int\frac{d\omega}{\omega}\alpha^{2}F(\omega),$ (10)
and can be used in the context of superconductivity to evaluate the critical temperature $T_{\rm c}$ assuming a phonon-mediated pairing mechanism and the Allen-Dynes formula [71]:
$T_{\rm c}^{\rm AD}=\frac{\hbar\omega_{\rm log}}{1.2k_{\rm B}}\mathrm{exp}{\left(\frac{-1.04(1+\lambda)}{\lambda-\mu^{*}(1+0.62\lambda)}\right)},$ (11)
where $k_{\rm B}$ is the Boltzmann constant, $\mu^{*}$ is the semi-empirical Coulomb potential, and $\omega_{\rm log}$ is the logarithmic average of the phonon frequency (geometric average) defined as
[…]
Although this formula is generally used for $s$-wave pairing mechanism with time reversal symmetry, a BCS type of formula can be obtained for all the other gap channels $p$, $d$, or $f$ with spin singlet and triplet type of pairing (breaking time reversal) [72, 73]. Although certain deviations are expected, one can still explore the possibility of this pairing mechanism in magnetic compounds by means of the Allen-Dynes formula.
Finally, we study resistive transport in magnetic metals. The transport of charge in a magnetic material depends on the spin-resolved carrier scattering rate $1/\tau_{n\mathbf{k}}^{\sigma}$ defined as
$\frac{1}{\tau_{n\mathbf{k}}^{\sigma}}\equiv\sum_{\mathbf{q}m\nu}w_{\mathbf{q}}|g_{mn\nu}^{\sigma}(\mathbf{k},\mathbf{q})|^{2}\\ \times[(n_{\mathbf{q\nu}}+1-f_{m\mathbf{k+q}}^{\sigma})\delta(\varepsilon_{n\mathbf{k}}^{\sigma}-\varepsilon_{m\mathbf{k+q}}^{\sigma}-\hbar\omega_{\mathbf{q}\nu})\\ +(n_{\mathbf{q\nu}}+f_{m\mathbf{k+q}}^{\sigma})\delta(\varepsilon_{n\mathbf{k}}^{\sigma}-\varepsilon_{m\mathbf{k+q}}^{\sigma}+\hbar\omega_{\mathbf{q}\nu})],$ (13)
where $f_{m\mathbf{k+q}}^{\sigma}$ is the Fermi-Dirac occupation for the spin $\sigma$ and $n_{\mathbf{q\nu}}$ is the Bose-Einstein distribution function. As for previous quantities, the average scattering rate is additive in the collinear approximation to magnetism:
$\langle{\tau}^{-1}\rangle\sim\sum_{\sigma}\sum_{n\mathbf{k}}w_{\mathbf{k}}\frac{1}{\tau_{n\mathbf{k}}^{\sigma}}.$ (14)
Prior studies have used the lowest-order variational approximation (LOVA) to study the resistivity of magnetic systems [49] which is based on computing the average scattering rate. This is a strong approximation in anisotropic systems or systems with different Fermi sheets [74]. Here, we solve the linearized Boltzmann transport equation (BTE) to calculate the conductivity [75, 76]
$\sigma_{\alpha\beta}^{\sigma}=\frac{-e}{V^{\rm uc}}\sum_{n\mathbf{k}}w_{\mathbf{k}}v^{\sigma}_{n\mathbf{k}\alpha}\partial_{E_{\beta}}f_{n\mathbf{k}}^{\sigma},$ (15)
where $V^{\rm uc}$ is the volume of the unit cell, $\partial_{E_{\beta}}f_{n\mathbf{k}}^{\sigma}$ is the perturbed occupation for an applied electric field $E_{\beta}$ in the direction $\beta$, $v_{n\mathbf{k}\alpha}^{\sigma}$ is the $\alpha$ component of the spin-resolved electronic band velocity, and $\sigma_{\alpha\beta}^{\sigma}$ is the conductivity tensor for spin index $\sigma$. The corresponding resistivity is
$\rho_{\alpha\beta}^{\rm BTE}\equiv(\sigma_{\alpha\beta}^{\uparrow}+\sigma_{\alpha\beta}^{\downarrow})^{-1}=\left[(\rho_{\alpha\beta}^{\uparrow})^{-1}+(\rho_{\alpha\beta}^{\downarrow})^{-1}\right]^{-1},$ (16)
which is the usual law of addition for resistivities in parallel, and in this case $\rho_{\alpha\beta}^{\sigma}$ is the resistivity experienced by the electrons in each spin channel. When comparing to polycrystalline samples, we use
$\rho^{\rm BTE}=\frac{\rm{Tr}[\rho_{\alpha\beta}^{\rm BTE}]}{3}.$ (17)
The out-of-equilibrium occupations $\partial_{E_{\beta}}f_{n\mathbf{k}}^{\sigma}$ [74] in the collinear magnetic case are obtained by solving the BTE
$\partial_{E_{\beta}}f_{n\mathbf{k}}^{\sigma}=ev_{n\mathbf{k}\beta}^{\sigma}\frac{\partial f^{\sigma}_{n\mathbf{k}}}{\partial\varepsilon^{\sigma}_{n\mathbf{k}}}\tau_{n\mathbf{k}}^{\sigma}+\frac{2\pi\tau_{n\mathbf{k}}^{\sigma}}{\hbar}\sum_{\mathbf{q}m\nu}w_{\mathbf{q}}|g_{mn\nu}^{\sigma}(\mathbf{k},\mathbf{q})|^{2}\\ \times[(n_{\mathbf{q\nu}}+1-f_{m\mathbf{k+q}}^{\sigma})\delta(\varepsilon_{n\mathbf{k}}^{\sigma}-\varepsilon_{m\mathbf{k+q}}^{\sigma}-\hbar\omega_{\mathbf{q}\nu})\\ \!+\!(n_{\mathbf{q\nu}}\!+\!f_{m\mathbf{k+q}}^{\sigma})\delta(\varepsilon_{n\mathbf{k}}^{\sigma}\!-\!\varepsilon_{m\mathbf{k+q}}^{\sigma}\!+\!\hbar\omega_{\mathbf{q}\nu})]\partial_{E_{\beta}}f_{m\mathbf{k+q}}^{\sigma}.$ (18)
Here $\partial f^{\sigma}_{n\mathbf{k}}/\partial\varepsilon^{\sigma}_{n\mathbf{k}}$ is the derivative of the Fermi-Dirac distribution. The equation is solved through an iterative procedure and if we neglect the second term in Eq. (18), we obtain the self-energy relaxation time approximation (SERTA) [75], which in the case of metallic systems is usually quite accurate.
III Methods
DFT calculations are performed using Quantum ESPRESSO (QE[60, 61] version 7.4 and Abinit version 10.2.5 [77, 56, 57, 59]. Scalar-relativistic norm-conserving pseudopotentials [78] from the standard accuracy table of the PseudoDojo project [79] are used. We use the local density approximation (LDA) [80] and the generalized gradient approximation (GGA) with the Pardew Burke Ernzerhof (PBE) functional [81]. The interpolated phonon dispersions and dynamical matrices are obtained with DFPT [42, 44, 43, 45]. The maximally localized Wannier functions are obtained using the wannier90 package version 3.1 [82]. Direct evaluation of spin-resolved EPME is obtained by modifying QE to print the values. We provide the modified code on the Materials Cloud Archive [83] along with all data to reproduce the results presented here. The interpolation of electronic band structures, phonon band structures, and electron phonon matrix elements, as well as the calculation of transport properties and electron-phonon coupling matrix elements are performed by modifying the EPW package version 5.9 [64]. This feature will be released in a future EPW version and we also provide the code on the Materials Cloud Archive [83].
The cutoff energy for the plane wave expansion, $\bf k$-meshes for the electrons, and the $\bf q$-meshes for the phonons are selected so that an arbitrary increase in those parameters produces a total energy change of less than 1 meV per unit cell. This corresponds to 100 Ry for both Fe and Ni, as well as a $\bf k$-point grid of 243 and a $\bf q$-point grid of 83 for both Ni and Fe. Structural relaxations are performed with a convergence criterion of $\Delta E<10^{-12}$ Ry, and with a pressure of less than 1 bar. In the case of phonons, the convergence criteria in order to obtain accurate force constants was tr2_ph $10^{-20}$ for QE while for Abinit the convergence criteria is on the energy $\Delta E<10^{-20}$ Ha. We converge the resistivity below 1% above 150 K using EPW, and validate the results with an in-house developed version of the ElectronPhonon.jl package [84, 85], which relies on the real-space EPME computed with the EPW code [63, 64] that we extend to the collinear magnetic case for the purpose of this work. Excellent agreement is obtained between the two codes, see Supplemental Material S1 [83] for more information.
IV Electron-phonon coupling with magnetism
In this work, we study BCC Fe and FCC Ni, both in their ferromagnetic configuration using the collinear approximation. We also consider non-magnetic Pb and verify that switching collinear magnetism produced the same result as non-magnetic calculations, as described in Supplementary Material S2 [83].
IV.1 Verification of the direct evaluation of the electron-phonon coupling
To verify our implementation, we compare the QE results with those obtained with Abinit. Although both codes match well in non-magnetic compounds [65, 66, 67], no validation study had yet been performed in the magnetic case. To this end, we compare basic quantities including total energy, total magnetization and phonon frequencies while keeping the computational parameters as close as possible. We find an excellent level of agreement between the two codes that is similar to the non-magnetic case, see Supplementary Material S3 [83].
Figure 1: Direct evaluation of ferromagnetic iron deformation potential for (a) spin up and (b) spin down at $\mathbf{k}=\Gamma$ for $m=n=5$ along a $\mathbf{q}$ momentum line for all phonon modes $\nu$. Red dots are computed with Quantum ESPRESSO (QE) while blue squares are computed with Abinit. The LDA functional is used.
We compare the absolute value of the deformation potentials between the two codes. To achieve this, we note that QE performs an average over all degenerate states in electronic and phonon band indices. We therefore applied the same average to the Abinit results and present a comparison for ferromagnetic Fe in Fig. 1 between the two codes for $\mathbf{k}=\bm{\Gamma}$ and bands $m=n=5$. We find a good agreement between the two codes with a maximum error of less than 0.02 meV/bohr. We provide the same comparison for Ni in Supplemental Material S4 [83] where we find an error below 0.01 meV/bohr.
IV.2 Verification of the electron phonon interpolation
Since we obtained reliable EPME by direct evaluation in QE, we use the EPW package to interpolate them into finer $\mathbf{q}$ and $\mathbf{k}$ grids. We extend the EPW package to be compatible with collinear magnetism and introduce two new input variables to control the selection of spin channel.
Figure 2: Electronic band structure for ferromagnetic (a) Fe and (b) Ni. Red/orange (blue/cyan) lines correspond to spin $\uparrow$ ($\downarrow$) or majority (minority) channel. Phonon band structure for (c) Fe and (d) Ni with LDA (green) and PBE (orange) functionals where the black dots are experimental values from Refs. [86, 87]. Solid (dashed) lines correspond to QE (EPW) calculations.
Figure 3: Deformation potential for the second to sixth lowest electronic bands in Fig. 2(a) and 2(b) for (a) Fe and (b) Ni at $\mathbf{k}=\bm{\Gamma}$ for $\uparrow$ (red) and $\downarrow$ (blue). Solid dots are direct evaluation of the deformation potential with Quantum ESPRESSO. The lines are the EPW interpolated deformation potential from $4^{3}$, $6^{3}$, $8^{3}$, $10^{3}$ and $12^{3}$ $\mathbf{k/q}$ homogeneous coarse grids with a color gradient from lighter to darker representing coarser to denser grids. Both spin channels are close in Ni while a larger difference is observed for Fe.
We first compare in Fig. 2(a) and 2(b) the direct and Wannier-interpolated electronic band structures of ferromagnetic Fe and Ni respectively, obtaining an energy mismatch of less than 0.1 meV for all the displayed $\mathbf{k}$-points and bands. This shows that both spin channels are correctly reproduced by the interpolation procedure. We then compute in Fig. 2(c) and 2(d) the phonon dispersion of ferromagnetic Fe and Ni respectively using LDA and PBE. Overall, LDA overestimates by about 10% the phonon frequency with respect to neutron scattering experiments [86, 87]. Using the PBE functional yields a much better agreement, in line with a previous study [88]. Moreover, we validate the interpolated electron-phonon deformation potential by comparing to direct calculations in Fig. 3, for electronic bands crossing the Fermi level and at the electronic zone center. We show the two spin channels for Fe and Ni, and find that convergence is achieved for $\times12$$\times$12 $\mathbf{k}/\mathbf{q}$ coarse grids in both cases. For such grids, we find excellent agreement between interpolated and independently calculated reference direct calculations. In addition, we also compare the EPW interpolation with the Abinit interpolation finding similar results; see Supplementary Material S4 [83]. Finally, we verify that the Hamiltonian, dynamical matrices, and EPME all decay rapidly in real space; see Supplementary Material S5 [83].
V Physical properties and electron-phonon quantities
Having validated our spin-resolved Wannier-interpolation scheme in magnetic materials, we compute physical observables linked with electron-phonon interactions. We start by computing the full electron-phonon coupling strength $\lambda$, and the corresponding spin-resolved $\lambda^{\sigma}$ using Eq. (7)
$\lambda^{\sigma}=\sum_{\mathbf{q}\nu}w_{\mathbf{q}}\lambda_{\mathbf{q}\nu}^{\sigma}$ (19)
and we report the results for ferromagnetic Fe and Ni in Table 1.
Compound Magnetism $\lambda$ $\lambda^{\uparrow}$ $\lambda^{\downarrow}$ $\omega_{\mathrm{log}}[\rm K]$ $T_{\rm c}^{\rm AD}[\rm K]$
Fe (BCC) FM 0.2961 0.2180 0.0781 311 0.02
Ni (FCC) FM 0.2767 0.0040 0.2727 291 0.01
Table 1: Total and spin-resolved electron phonon coupling strength $\lambda$ for ferromagnetic (FM) Fe and Ni. Evaluated characteristic frequency $\omega_{\mathrm{log}}$ and Allen-Dynes critical temperature $T_{\mathrm{c}}^{\mathrm{AD}}$ with $\mu^{*}=0.13$.
Interestingly, we find that the calculated spin-resolved electron phonon coupling strength shows large differences between the two spin channels in both Fe and Ni. Although this behavior is not entirely surprising for Fe given the clear difference in deformation potential between the two spin channels, the same cannot be said in the case of Ni, whose splitting in deformation potential is much weaker, likely due to its smaller exchange splitting. Additionally, we note that the main contribution to the electron phonon coupling strength in the case of Fe comes from the spin up (majority) channel, while in the case of Ni we find the reverse situation. In order to understand this behavior, we inspect the definition of the spin-resolved electron-phonon coupling strength Eqs. (7) and (19) and we notice two main sources for this discrepancy. The first one is the deformation potential. The second is the electronic nesting for the given spin channel provided by the product of the two delta functions. This quantity is related to the density of states at the Fermi level. As seen in Fig. 3, the deformation potential is very different between the two spin channels for both Fe and Ni. In the case of Fe the difference in the deformation potential is significant and could explain the difference. However, $\lambda^{\downarrow}$ is smaller than $\lambda^{\uparrow}$, while the trend in the deformation potential is opposite, see Fig. 3(a). Consequently, the deformation potential cannot be the source of the difference between $\lambda^{\uparrow}$ and $\lambda^{\downarrow}$. In the case of Ni, both spin channels have similar values in the deformation potential, while the electron phonon coupling strength $\lambda$ is completely dominated by the spin down channel $\lambda\approx\lambda^{\downarrow}$. Therefore, the deformation potential is not the main driver for such differences in either of the two considered cases.
Figure 4: Spin resolved density of states for ferromagnetic (a) Fe and (b) Ni. Positive and negative density of states representing up (red) and down (blue) spin channels.
Inspecting the density of states close to the Fermi level for each compound in Fig. 4, we note that for Fe the spin up channel (majority) is 2.3 times larger than the spin down channel (minority), see Fig. 4(a), while the opposite behavior is found for Ni where the spin down (minority) channel is 7.6 times larger, see Fig. 4(b). This is a well-known effect linked to the d-band filling. Thus, the differences in the spin-resolved electron-phonon coupling strength $\lambda^{\sigma}$ between the two spin channels for Fe and Ni are mainly related to Fermi-level nesting and the density of states at the Fermi level. Our results for the electron-phonon coupling strength of Fe compare well with those obtained in Ref. [70].
Figure 5: Eliashberg spectral function for ferromagnetic (a) Fe and (b) Ni. Red lines corresponds to the spin up (majority) and blue line corresponds to spin down (minority).
We also compute the Eliashberg spectral functions and show the result in Fig. 5. We see in Fig. 5(a) that the Fe spin majority channel is the dominant contribution for most of the energy range. In contrast, for Ni the spin minority dominates throughout and the spin majority is negligible, see Fig.  5(b). This important difference is due to the density of states at the fermi level in Ni, see Fig. 4. Using the $\alpha^{2}F^{\sigma}(\omega)$, we evaluate the $\omega_{\mathrm{log}}$ and compute the $T_{\rm c}$ with the Allen-Dynes formula. In doing so, we need to provide a reasonable value for the semi-empirical Coulomb potential $\mu^{*}$. Typical values range between 0.1 and 0.2 in the literature  [89, 90, 91] and we choose $\mu^{*}=0.13$ here. We find critical temperatures of 0.02 K and 0.01 K respectively for Fe and Ni (see Table 1), showing the absence of Eliashberg superconductivity even if the pairing would not be affected by the internal magnetization. Performing a scan from $\mu^{*}=0.1-0.2$ we find that in the case of Fe, the $T_{c}$ ranges from 0.13 to 0 K while in the case of Ni $T_{c}$ ranges from 0.06 to 0 K. We note that magnetism could lead to a larger effective $\mu^{*}$ since it needs to account for the Coulomb potential and spin fluctuations [89], which would further decrease $T_{\rm c}$, even within Eliashberg theory.
The results obtained on the $T_{\rm c}$ of Fe and Ni are compatible with ambient pressure experiments, which never show a superconducting transition. Fe may show a superconducting transition at very low temperature $T_{c}\leq 2$ K under high pressure [92, 93].
V.1 Resistivity of magnetic metals
We now focus on the resistivity of metals and solve the BTE for both spin channels in ferromagnetic Fe and Ni. We note that when converging the resistivity and in general electron-phonon related quantities, it is possible to use different $\mathbf{k}/\mathbf{q}$ coarse and fine grids for the two spin channels since there is no guarantee that both spin channels would behave equivalently. Nevertheless, in our case, we find that convergence is achieved with the same settings. A good compromise between accuracy and computational cost is obtained with a $8^{3}$ $\mathbf{k/q}$ coarse grid; see Supplementary Material S7 [83] for further discussion.
The computed resistivity of Fe is presented in Fig. 6(a) and shows that the BTE with electron-phonon scattering explains only partially the resistivity and tends to underestimate it. We find that PBE results are slightly closer to experiment due to the improved phonon dispersion, see Fig. 2(c), and a small decrease in Fermi velocity; see Supplementary material S8 [83]. We also find that the SERTA is a good approximation to the iterative BTE solutions, which is typical for non-magnetic metals [94]. Our results are qualitatively similar to a previous study using the Ziman formula [49] where quantitative agreement with the experimental results is obtained at low temperatures. However, we notice that the electron-phonon contribution to resistivity does seem to dominate at low temperature, see inset in Fig 6(a), and at temperatures up to 300 K our calculations suggest that electron-phonon scattering can account for at least 75% of the experimental resistivity when using the PBE functional. For higher temperatures, the observed underestimation indicates that electron-phonon scattering is not the sole mechanism contributing to the resistivity.
Figure 6: Electrical resistivity as a function of temperature for (a) Fe and (b) Ni, with an inset at low temperature 0-200 K. Black symbols are obtained from Refs. [95, 96, 97]. Lines are computed resistivities using SERTA (dashed) and full BTE (solid) with LDA (green) and PBE (orange).
We continue our analysis by investigating the case of ferromagnetic Ni. As seen in Fig. 6(b), the computed resistivity for Ni is also underestimated in the entire temperature range compared to experiment. However, above 350 K the resistivity originating from electron-phonon scattering is almost one order of magnitude less than the experimental resistivity. In addition, the agreement at low temperatures is worst than in Fe since there is an important underestimation of the resistivity already in the 50-100 K temperature range, see the inset in Fig. 6(b). Interestingly, we note that using the PBE functional does not improve significantly in spite of the remarkably better agreement of the phonon dispersion with experiments, see Fig.2(d). This hints at an important distinction with respect to Fe.
In ferromagnetic compounds, contributions to the resistivity originating from magnons or spin fluctuations are expected to be significant [98, 99]. One of the reasons why magnons may explain the missing resistivity of Ni and Fe above 300 K is due to their energy scale which is around 8 to 9 times bigger than the phonons [100]. The magnon population scales as $(T/T_{\rm Curie})^{2}$ [99], and will be smaller in Fe than in Ni for a given temperature due to the lower Curie temperature ($T_{\rm Curie}$) of 627 K  [96], compared to 1043 K in Fe [97]. Consequently, the saturation that gives a linear relationship with the phonons occurs at significantly elevated temperatures, and the “low temperature” $T^{2}$ behavior of the magnons is observed over a larger range of temperatures. Although this explains the increase in the importance of magnon at high temperature, it does not explain the marked difference between Fe and Ni.
In fact, the magnon population is crucially dependent on the density of states at the fermi level of the opposite spin channel [101]. The large density of states of the spin down (minority) channel in Ni (see Fig. 4(b)) above the Fermi level allows for a large magnon creation by converting spin up (majority) electrons into spin down (minority) electrons. In the case of Fe, magnon creation is suppressed by the small accessible phase space of minority electrons [101]. In addition to the difference in magnon creation, our calculations show that the spin up electrons in Ni experience far less scattering from electron-phonon, and, as a result, our calculations show that $\rho_{\alpha\beta}\approx\rho^{\uparrow}_{\alpha\beta}$ with an order of magnitude difference in the conductivities $\sigma_{\alpha\beta}^{\uparrow}\gg\sigma^{\downarrow}_{\alpha\beta}$, while for Fe $\rho^{\uparrow}_{\alpha\beta}\sim\rho_{\alpha\beta}^{\downarrow}$. Since the spin up channel for Ni shows a larger lifetime broadening due to magnons [101], the total lifetime $\tau_{n\mathbf{k}}^{\uparrow}$ is smaller than with electron-phonon scattering only, explaining the underestimated resistivity compared to experiment in Ni.
In summary, comparing the calculated resistivity at low temperature with the experimental data of Fe and Ni, we find that the behavior of these two compounds is fundamentally different: magnons dominate the scattering in Ni while electron-phonon is the dominant scattering mechanism in Fe up to room temperature.
V.2 Role of magnetism on physical properties
Figure 7: Phonon dispersion for spin-unpolarized (non-magnetic) (a) Fe and (b) Ni obtained with PBE functional. Black dots are the experimental data from Refs. [86, 87].
One of the key aspects of including magnetism is the fact that the electronic structure of the compounds changes with magnetism, as does the phonon dispersion. As seen in Fig. 7(a), Fe is not dynamically stable when we use the non-magnetic (NM) approximation. This is due to a fermionic pressure that increases the lattice parameters of Fe when magnetism is allowed [102, 103, 104]. The instabilities found in the phonon dispersion for a BCC structure in the NM approximation are known to be related to distortions that lead to hexagonal phases [105]. This is also consistent with the fact that in the crystallographic phase diagram of Fe, we find a hexagonal phase at pressures a bit above 10 GPa [106, 107]. It is also reported that the FM phase becomes paramagnetic, antiferromagnetic or even diamagnetic at sufficiently high pressures.
Interestingly, the phonon dispersion for Ni does not change much by using the NM approximation; see Fig. 7(b). In fact, one can notice that the FM phase of Ni produces only a small overall imbalance in the spin occupations since the experimental magnetic moment per unit cell is $M_{\mathrm{exp}}=0.616~\mu_{\mathrm{B}}$ [108] when compared with Fe that presents a magnetic moment of $M_{\mathrm{exp}}=(2.2\pm 0.1)~\mu_{\mathrm{B}}$ [109, 110, 111]. Thus, one may wonder how different the resistivity in the NM phase is, when compared with that in the FM collinear case for Ni. In that regard, we repeat the calculations with the BTE for the resistivity of Ni with PBE (see LDA in the supplementary material S10 [83]) but using the NM phase and we compare it in Fig. 8(b) with the FM case and with the experimental data.
Figure 8: Electrical resistivity of iron and nickel as a function of temperature. The solid green and purple lines are the computed PBE resistivities $\rho^{\rm BTE}$ using ferromagnetic and non-magnetic EPC, respectively. Black dots are experimental values obtained from Refs. [95, 96, 97].
We find that the NM approximation produces a slight overestimation of the resistivity for temperatures below 300 K and then it underestimates the resistivity.
Although the limitation of the NM approximation has been highlighted previously [112, 113, 114, 115, 116], it is common to find studies that still use the NM approximation to model spin-disordered paramagnetic, the ferromagnetic or antiferromagnetic phases. For this reason, we also compute the resistivity of NM Fe as a function of temperature and present it in Fig. 8(a). As expected, we obtain an extremely overestimated resistivity when compared to the FM one since we neglect scattering originating from soft modes. Importantly, we show that even though the phonons are similar in Ni, neglecting the spin degree of freedom yields the wrong conclusion that the resistivity is dominated by electron-phonon interactions. We remark that, in spite of the seemingly better agreement with the experiment, the NM results actually overestimate the experimental data, which should never happen since we are not considering the full set of scattering channels. Moreover, the temperature-dependent trend of the NM result is inconsistent with experimental observation.
VI Conclusions
We explore the role of electron-phonon interactions in typical magnetic metals. Namely ferromagnetic Fe and Ni where we show that neither the deformation potential nor the electron-phonon coupling present the same strength for the two spin channels. The latter being dominated by Fermi surface nesting effects and the partial density of states of each spin channel. Additionally, we show that the electrical resistivity of Fe and Ni presents important contributions from electron-phonon scattering at room temperature although the relative importance is material dependent. While in Fe the resistivity due to electron-phonon interactions is the dominant contribution at room temperature, for Ni it is less than one-third of the total. Additionally, we show that including the spin degree of freedom to compute resistivity is crucial. Neglecting it can lead to unstable phonons such as in the case of Fe, can overestimate resistivity, and can yield a completely erroneous interpretation for the source of the resistivity. Furthermore, we explore the phonon-mediated superconducting mechanism in these two elemental compounds, a subject rarely addressed with proper numerical calculations. We find suppressed superconductivity, showing that even in a non-magnetic configuration their critical temperatures would be extremely low. These results validate and extend the capabilities of EPW to compute electron-phonon quantities with collinear magnetism. Beyond the methodological advances, our findings highlight how magnetism critically shapes resistive losses and superconductivity in metals that are central to energy applications. By enabling accurate modeling of transport in magnetic systems, this work provides a foundation for the design of low-energy-consumption spintronic devices, improved conductive pathways with reduced losses, and future magnetic materials for energy conversion and transmission. Future extensions to include spin–orbit coupling and non-collinear magnetism will further expand the reach of this framework, opening new directions for first-principles studies of complex magnetic materials.
Acknowledgements.
S. P. is a Research Associate and M.M. is an Aspirant PhD candidate of the Fonds de la Recherche Scientifique - FNRS. This work was supported by the Fonds de la Recherche Scientifique - FNRS under Grants number T.0183.23 (PDR). M.J.V. and G.A. acknowledge the Fonds de la Recherche Scientifique (FRS-FNRS Belgium) and Fonds Wetenschappelijk Onderzoek (FWO Belgium) for EOS project CONNECT (G.A. 40007563), and Fédération Wallonie Bruxelles and ULiege for funding ARC project DREAMS (G.A. 21/25-11). M.J.V. acknowledges funding by the Dutch Gravitation program “Materials for the Quantum Age” (QuMat, reg number 024.005.006), financed by the Dutch Ministry of Education, Culture and Science (OCW). Computational resources have been provided by the supercomputing facilities of the Université catholique de Louvain (CISM/UCL) and the Consortium des Équipements de Calcul Intensif en Fédération Wallonie Bruxelles (CÉCI) funded by the Fond de la Recherche Scientifique de Belgique (F.R.S.-FNRS) under convention 2.5020.11 and by the Walloon Region. The present research benefited from additional computational resources made available on Lucia, the Tier-1 supercomputer of the Walloon Region, infrastructure funded by the Walloon Region under the grant agreement n°1910247, and by EuroHPC (Extreme grant EHPC-EXT-2023E02-050) on Marenostrum5 at BSC, Spain.